# Spectral Flow Learning Theory: Finite-Sample Guarantees for Vector-Field Identification

### v3

@@ III-A General Form
Let there be a context window of length $M$ with step-window $h_{1},\dots,h_{M-1}]\in\boldsymbol{h}$. Denote the future step as the window anchor, $h\coloneqq h_{M}$. The bounded step-ratio vector is defined as:
$\zeta\coloneqq[\zeta_{1},\dots,\zeta_{M-1}]\in\boldsymbol{\zeta}\subset(0,\bar{\zeta}]^{M-1},$
where $\zeta_{k}\coloneqq h_{k}/h_{k-1}$ and $\zeta_{M-1}=h/h_{M-1}$.
@@
A vLMM takes the following general form [5, Eq. 5.15]:
@@
with coefficients: $\alpha_{jk}\coloneqq\alpha_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1}),\beta_{jk}\coloneqq\beta_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1}),$ functions of the step ratios $\zeta$.
Before stating the deterministic bias assumptions, we instantiate the unified form (3) with the standard variable-step families: Adams-Bashforth (AB), Adams-Moulton (AM), and backward-differentiation formulas (BDF) [5]. Each of the above method arises from a suitable choice of the ratio-dependent coefficients, $(\alpha_{jk}(\zeta),\beta_{jk}(\zeta))$, and admits a windowed-flow representation aligned with our SFL label operator, bringing explicit, implicit, and stiff-stable linear discretizations under one single framework. Furthermore, notice that the popular Savitzky-Golay (SG) derivative estimators arise from local polynomial least-squares fitting on a sliding window [6, 7]. Therefore, for uniform steps and one-sided windows, SG estimators reduce to the classical finite-difference derivative weights underlying multistep formulas. Within SFL framework, SG estimators can be seen as an alternative windowed label operator providing derivative labels complementary to vLMM flow labels.
@@ III-B Assumptions for Deterministic Bias
A vLMM (3) is order-$p$  [5, Def. 5.2] if:
$q(x_{k_{M}})+\sum_{j=1}^{M-1}\alpha_{jk}q(x_{k+j})=h_{k+M-1}\sum_{j=0}^{M}\beta_{jk}q^{\prime}(x_{k+j})$
holds for all polynomials $q(x)$ of degree $\leq p$ and grids $(h_{j})$.
Notice that the window size $M$ is the number of back steps a $M$-step vLMM uses, whereas the order $p$ is the accuracy of the method. They are related but not identical: in classical families, AB has $p=M$, AM has $p=M+1$, and BDF has $p=M$ (with zero-stability only up to $M\leq 6$). Moving onward, we state the key assumptions needed from the vLMM theory to establish an FS-HP bound in the SFL framework.
@@
$\underline{\zeta}\ \leq\ \zeta_{k}\ \leq\ \bar{\zeta}\qquad\forall k.$
Moreover, the variable–step coefficients $\{\alpha_{jk},\beta_{jk}\}$ are uniformly bounded.
Notice that bounded coefficients $\alpha_{jk},\beta_{jk}$ follow from bounded step ratios for AB/AM methods, and BDF [5, Lem. 5.3].
@@
The right–hand side $f$ is sufficiently smooth, i.e., $f\in\mathcal{C}^{p+1}$ and locally Lipschitz on the forward–invariant compact set $\mathcal{X}$.
@@ IV-A Population and Sampling Hilbert Spaces
where $\mathcal{H}_{\mathcal{V}}$ is a vvRKHS equipped with an inner product, $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathcal{V}}}$, induced by the measurable, bounded PHS kernel $k_{\rm phs}$ and densely embedded in $\mathcal{V}$ with the inclusion map $\mathsf{I}_{\mathcal{V}}:\mathcal{H}_{\mathcal{V}}\to\mathcal{V}$. Recall that $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}\coloneqq\rho_{\mathcal{X}}^{\mathrm{design}}\times\rho_{\boldsymbol{h}}$, the single-step flow space is:
@@
with the embedded vvRKHS $\mathcal{H}_{\mathcal{W}^{\circ}}$ and the inclusion map defined as $\mathsf{I}_{\mathcal{W}^{\circ}}:\mathcal{H}_{\mathcal{W}^{\circ}}\to\mathcal{W}^{\circ}$. Notice that $\mathcal{W}^{\circ}$ is the space in which the one-step exact flow $\Phi^{h}(x;f_{\rho})$ lives. For the windowed flow space, recall that the window anchor is $h_{\rm anc}\in h$ and the bounded step-ratio vector is:
$\zeta\coloneqq[\zeta_{1},\dots,\zeta_{M-1}]\in\boldsymbol{\zeta}\subset(0,\bar{\zeta}]^{M-1},$
where $\zeta_{k}\coloneqq h_{k}/h_{k-1}$ and $\zeta_{M-1}=h_{\rm anc}/h_{M-1}$. Denote the backshift times by $\tau_{j}(\zeta)\geq 0$ measured from $h_{\rm anc}$ and a composite measure $\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}\coloneqq\rho_{\mathcal{X}}^{\rm design}\times\rho_{\boldsymbol{h}}\times\rho_{\boldsymbol{\zeta}}$. Define the windowed flow space:
$\mathcal{W}\coloneqq L^{2}\big(\mathcal{X}\times\boldsymbol{h}\times\boldsymbol{\zeta},\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}};\ \mathbb{R}^{n}\big),\quad\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}.$
This is the space in which spectral regularization acts directly.
@@
$(\mathsf{J}\Phi)(x,h,\zeta)\coloneqq\Phi^{h}(x),$
@@
$\Phi^{h}(x;f)=(\mathsf{J}\Phi)(x,h,\zeta;f)\eqqcolon\Psi(x,h,\zeta;f)\in\mathcal{W}.$
@@ IV-B Inclusion Maps and Other Important Operators
We use spectral‑filter estimators $\Phi^{\lambda_{\ell}}_{z,y}\in\mathcal{H}_{\mathcal{W}}$ based on a filter ${\mathsf{g}}_{\lambda}$ applied to $\mathsf{T}_{\mathcal{W}}$ (vector‑valued extension is componentwise).
@@ IV-C Filter Qualification and Lipschitz Condition
Another component we need for constructing a FS-HP bound is the notion of a spectral filter, its qualification, and the Lipschitz condition.
@@
[4, Def. 1] Let ${\mathsf{g}}_{\lambda}:[0,\kappa^{2}]\to\mathbb{R}$ be a spectral filter. Its qualification $\nu>0$ is the largest order such that:
@@
A spectral filter is an approximate inverse applied to the spectrum of a positive self-adjoint operator (e.g., the covariance/integral operator) via functional calculus: if a covariance operator $\mathsf{T}=\sum_{i}\sigma_{i}\langle\cdot,e_{i}\rangle e_{i}$, then ${\mathsf{g}}_{\lambda}(\mathsf{T})=\sum_{i}{\mathsf{g}}_{\lambda}(\sigma_{i})\langle\cdot,e_{i}\rangle e_{i}$. It attenuates directions with small eigenvalues $\sigma_{i}$ to stabilize the inversion, while the residual $1-\sigma{\mathsf{g}}_{\lambda}(\sigma)$ quantifies the bias introduced by regularization. The qualification $\nu$ is the largest order for which $r_{\lambda}(\sigma)\leq\lambda^{\nu}\sigma^{-\nu}$ uniformly. Typical values of $\nu$ are: Tikhonov with $\nu=1$; order-$t$ iterated Tikhonov with $\nu=t$; Landweber with $\nu\to\infty$ [4, Sec. 3.1].
Consequently, under a source condition of order $r\leq\nu$ one obtains approximation error of order $\lambda^{2r}$ (in the prediction norm controlled by $\mathsf{T}$). The Lipschitz condition ensures stability of ${\mathsf{g}}_{\lambda}(\mathsf{T})$ to empirical perturbations $\mathsf{T}\mapsto\mathsf{S}_{\ell}\circ\mathsf{T}$, which is crucial for finite-sample FS-HP bounds. Gaussian process regression can be seen as Tikhonov regularization with $\lambda=\sigma_{y}^{2}/\ell$, with $\sigma_{y}^{2}$ the variance of observation noise.
@@ IV-D Source Condition
In the previous section, we fix the hypothesis spaces $(\mathcal{H}_{\mathcal{V}},\mathcal{H}_{\mathcal{W}})$ and the spectral filter ${\mathsf{g}}_{\lambda}$, with qualification $\nu$ and Lipschitz exponent $\mu$. Moving onward, we quantify target-kernel alignment by a Hölder-type source condition [1] with exponent $r>0$; larger $r$ yields faster rates up to the filter’s qualification $\nu$. Conceptually, the source condition measures how well the targets $\Phi_{\rho}$ and $f_{\rho}$ align with the vvRKHS prior; formally, it requires them to lie in the range of a positive power of the integral operator $\mathsf{L}=\mathsf{I}\mathsf{I}^{*}$, and its vector-field analogue.
SFL learns dynamics in a windowed flow space and lifts both estimates and their guarantees to the vector field. Accordingly, we impose source conditions at both levels: a flow condition that directly drives the flow FS-HP bound, and a vector-field condition that encodes differential physics, e.g., Hamiltonian passivity, incompressibility $\nabla\cdot f=0$, and symmetry/equivariance. In practice, these constraints can be enforced via a physics-informed vvRKHS kernel [8], by projecting the lifted estimator onto the constraint set [9], or by adding a physics-informed penalty to the loss [10].
For $r>0$ and $R_{\mathcal{V}}>0$, assume $f_{\rho}$ is the target vector field with the following structure defined through (4):
@@
where $\mathsf{L}_{\mathcal{W}}$ is define in (5). This flow-level source is the one used in Theorem 1 to obtain the FS-HP rate $\ell^{-\frac{2r}{2r+\beta}}$ (with $\beta=\max\{1,2\mu\}$).
Remark IV.1 (Compatibility/transferring source between levels).
If one wishes to derive the flow source from the vector-field source, it suffices to choose $\mathcal{H}_{\mathcal{W}}$ compatible with the exact flow labeling map $\mathsf{E}_{\rm exact}:\mathcal{H}_{\mathcal{V}}\to\mathcal{W}^{\circ}$ and the lifting $\mathsf{J}:\mathcal{W}^{\circ}\to\mathcal{W}$. Here $\mathsf{E}_{\rm exact}$ is the exact one-step flow map that sends a vector field to its true flow: $(\mathsf{E}_{\rm exact}f_{\rho})(x,h)\coloneqq\Phi^{h}(x;f_{\rho})$, and composing with $\mathsf{J}$ yields the $\zeta$-constant embedding $(\mathsf{J}\mathsf{E}_{\rm exact}f_{\rho})(x,h,\zeta)=\Phi^{h}(x;f_{\rho})$. Under this design, there is a constant $C_{\rm push}$ (depending only on operator norms) such that:
$f_{\rho}\in\Omega^{(\mathcal{V})}_{r,R_{\mathcal{V}}}\quad\Longrightarrow\quad\Phi_{\rho}=\mathsf{J}\mathsf{E}_{\rm exact}f_{\rho}\ \in\ \Omega^{(\mathcal{W})}_{r,R_{\mathcal{W}}}$
with $R_{\mathcal{W}}\leq C_{\rm push}R_{\mathcal{V}}$. In our general theory, we assume the flow source directly (to remain agnostic about the precise hypothesis design in $\mathcal{W}$), but the construction above shows how to enforce it by design.
@@ V Spectral Flow Learning Theory
This section states the core constructions and guarantees of SFL. We formalize how variable-step multistep structure induces well-posed labels via the Koopman-lag forcing and state-aggregation operators, define the label map and vLMM residual, and provide an observability inequality (with constant $c_{\rm obs}$) together with a uniform local-truncation bound. These ingredients combine with spectral regularization in the windowed flow space to yield FS-HP-style, excess-risk bounds in the flow space and, via observability, in the vector-field space.
@@ V-A Label Map and vLMM Residual
A key aspect of a theory on learning a vector field from trajectory data is the construction of well-defined input-label pairs. In our framework, we leverage the Koopman-lag forcing operator and state-aggregation operator to define the population level loss through the label map and vLMM residual.
Consider a vLMM as defined in Definition 1. Fix the window and anchor it at $t_{k+M}$. Define the backward lags:
@@
Using the Koopman action $(U^{t}f)(x)=f(\Phi^{t}(x))$, the right-hand side of (3) can be rewritten as:
$h_{k+M-1}\sum_{j=1}^{M}\beta_{jk}(U^{-\tau_{M-j}}f)\big(x_{k+M}\big).$
@@
$(\mathsf{B}f)(x;h,\zeta)\coloneqq h\sum_{j=0}^{M}\beta_{j}(\zeta)(U^{-\tau_{j}}f)(x)\quad:\ \mathcal{V}\to\mathcal{W},$ (6)
where $\tau_{j}(\zeta)$ are the cumulative lags determined by $\zeta$ and the anchor step $h\coloneqq h_{k+M-1}$. Similarly, the state-aggregation operator is defined as:
$(\mathsf{A}\Phi)(x;h,\zeta)\coloneqq\sum_{j=1}^{M}\alpha_{j}(\zeta)\Phi^{-\tau_{j}}(x)\quad:\ \mathcal{W}^{\circ}\to\mathcal{W}.$
Now, we are ready to define the population level label map and vLMM residual.
@@
The population label map and vLMM residual are:
$\mathsf{Y}_{\alpha}(\Phi)\coloneqq\mathsf{J}\Phi+\mathsf{A}\Phi\in\mathcal{W},\ \Delta_{\mathrm{MS}}(f)\coloneqq\mathsf{Y}_{\alpha}(\Phi)-\mathsf{B}f\ \in\mathcal{W},$ (7)

### v4

@@ III-A General Form
The vLMM general form provides a linear map between an ODE and its solution on an irregular time grid $h_{j}]_{-\infty}^{\infty}$. Recall that a step-window of length $M$ on a time grid is defined as $T_{M}\coloneqq(h_{\rm anc},\zeta(M-1))$. For each solution $x_{k+M}$, we consider a step-window $T_{M}(k)=(h_{k+M-1},\zeta_{k}(M-1))$, with $\zeta_{k}(M-1)\coloneqq[\zeta_{k},\dots,\zeta_{k+M-1}]$.
@@
A vLMM takes the following general form [8, Eq. 5.15]:
@@
with coefficients: $\alpha_{jk}\coloneqq\alpha_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1}),\beta_{jk}\coloneqq\beta_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1})$, functions of the step ratios $\zeta_{k}(M-1)$.
Before stating the deterministic bias assumptions, we instantiate the unified form (3) with the standard variable-step families: Adams-Bashforth (AB), Adams-Moulton (AM), and backward-differentiation formulas (BDF) [8]. Each of the above methods arises from a suitable choice of the ratio-dependent coefficients, $(\alpha_{jk}(\zeta),\beta_{jk}(\zeta))$, and admits a windowed-flow representation aligned with our SFL label operator, bringing explicit, implicit, and stiff-stable linear discretizations under one single framework. Furthermore, notice that the popular Savitzky-Golay (SG) derivative estimators arise from local polynomial least-squares fitting on a sliding window [12, 13]. Therefore, for uniform steps and one-sided windows, SG estimators reduce to the classical finite-difference derivative weights underlying multistep formulas. Within SFL framework, SG estimators can be seen as an alternative windowed label operator providing derivative labels complementary to vLMM flow labels.
@@ III-B Assumptions for Deterministic Bias
[8, Def. 5.2] A vLMM (3) is order-$p$ if:
$q(x_{k+M})+\sum_{j=1}^{M-1}\alpha_{jk}q(x_{k+j})=h_{k+M-1}\sum_{j=0}^{M}\beta_{jk}q^{\prime}(x_{k+j})$
holds for all polynomials $q(x)$ of degree $\leq p$ and grids $h_{j}]_{-\infty}^{\infty}$.
Notice that the window size $M$ is the number of back steps a $M$-step vLMM uses, whereas the order $p$ is the accuracy of the method. They are related but not identical: in classical families, AB has $p=M$, AM has $p=M+1$, and BDF has $p=M$, with zero-stability only up to $M\leq 6$. Moving onward, we state the key assumptions needed from the vLMM theory to establish an FS-HP bound in the SFL framework.
@@
$\underline{\zeta}\ \leq\ \zeta_{k}\ \leq\ \bar{\zeta},\qquad\forall k.$
Moreover, the variable-step coefficients $\{\alpha_{jk},\beta_{jk}\}$ are uniformly bounded.
Notice that bounded coefficients $\alpha_{jk},\beta_{jk}$ follow from bounded step ratios for AB/AM methods and BDF [8, Lem. 5.3].
@@
The right-hand side $f$ is sufficiently smooth, i.e., $f\in\mathcal{C}^{p+1}$ and locally Lipschitz on the forward-invariant compact set $\mathcal{X}$.
@@ IV-A Population and Sampling Hilbert Spaces
where $\mathcal{H}_{\mathcal{V}}$ is a vvRKHS equipped with an inner product, $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathcal{V}}}$, induced by the measurable, bounded port-Hamiltonian kernel $k_{\rm phs}$ [3] and densely embedded in $\mathcal{V}$ with the inclusion map $\mathsf{I}_{\mathcal{V}}:\mathcal{H}_{\mathcal{V}}\to\mathcal{V}$. Recall that $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}\coloneqq\rho_{\mathcal{X}}^{\mathrm{design}}\times\rho_{\boldsymbol{h}}$, the single-step flow space is:
@@
with the embedded vvRKHS $\mathcal{H}_{\mathcal{W}^{\circ}}$ and the inclusion map defined as $\mathsf{I}_{\mathcal{W}^{\circ}}:\mathcal{H}_{\mathcal{W}^{\circ}}\to\mathcal{W}^{\circ}$. Notice that $\mathcal{W}^{\circ}$ is the space in which the one-step exact flow $\Phi^{h}(x;f_{\rho})$ lives. For the windowed flow space, we recall the $M$-steps window $T_{M}\in\mathcal{T}$ and its distribution $\rho_{\mathcal{T}}$. Denote the backshift times by $\tau_{j}(\zeta)\geq 0$ measured from $h_{k+M-1}$ and a composite measure $\rho^{\rm design}_{\mathcal{X},\mathcal{T}}\coloneqq\rho_{\mathcal{X}}^{\rm design}\times\rho_{\mathcal{T}}$. Define the windowed flow space:
$\mathcal{W}\coloneqq L^{2}\big(\mathcal{X}\times\mathcal{T},\rho^{\rm design}_{\mathcal{X},\mathcal{T}};\ \mathbb{R}^{n}\big),\quad\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}.$
The Hilbert space $\mathcal{W}$ is the space in which spectral regularization acts directly.
Remark IV.1.
@@
$(\mathsf{J}\Phi)(x;h,\zeta)\coloneqq\Phi^{h}(x),$
@@
$\Phi^{h}(x;f)=(\mathsf{J}\Phi)(x;h,\zeta,f)\eqqcolon\Psi(x,T;f)\in\mathcal{W}.$
@@ IV-B Inclusion Maps and Other Important Operators
In spectral regularization learning theory, we view the finite dataset as inducing empirical analogues of these operators. Given a sample $\mathcal{D}=\{(z_{i},y_{i})\}_{i=1}^{\ell}$ with windowed inputs $z_{i}:=(x_{i},T_{M}(i))$ and outputs $y_{i}\in\mathbb{R}^{n}$, let $\mathsf{S}_{z}:\mathcal{H}_{\mathcal{W}}\to\mathcal{Y}$ be the sampling operator $(\mathsf{S}_{z}\Phi)_{i}\coloneqq\Phi(z_{i})$ with empirical norm $\|v\|_{\mathcal{Y}}^{2}$. The empirical covariance and empirical integral operators are:
$\widehat{\mathsf{T}}_{\mathcal{W}}\coloneqq\mathsf{S}_{z}^{*}\mathsf{S}_{z}:\mathcal{H}_{\mathcal{W}}\to\mathcal{H}_{\mathcal{W}},\qquad\widehat{\mathsf{L}}_{\mathcal{W}}\coloneqq\mathsf{S}_{z}\mathsf{S}_{z}^{*}:\mathcal{Y}\to\mathcal{Y}.$
In coordinates, $\widehat{\mathsf{T}}_{\mathcal{W}}$ corresponds to the normalized information matrix built from feature vectors, while $\widehat{\mathsf{L}}_{\mathcal{W}}$ reduces to the kernel Gram matrix acting on sample outputs. Spectral regularization theory views $\widehat{\mathsf{T}}_{\mathcal{W}}$ and $\widehat{\mathsf{L}}_{\mathcal{W}}$ as finite-rank perturbations of $\mathsf{T}_{\mathcal{W}}$ and $\mathsf{L}_{\mathcal{W}}$, so that learning becomes an inverse problem of recovering a target function through nearly singular operators. The spectral filter $\mathsf{g}_{\lambda}$ stabilizes this inversion by damping directions associated with small eigenvalues. We consider estimators $\Phi^{\lambda_{\ell}}_{z,y}\in\mathcal{H}_{\mathcal{W}}$ with spectral filter ${\mathsf{g}}_{\lambda}$ applied to $\mathsf{T}_{\mathcal{W}}$.
@@ IV-C Filter Qualification and Lipschitz Condition
Another component we need for constructing the FS-HP bound is the notion of a spectral filter, its qualification, and the Lipschitz condition.
@@
[11, Def. 1] Let ${\mathsf{g}}_{\lambda}:[0,\kappa^{2}]\to\mathbb{R}$ be a spectral filter. Its qualification $\nu>0$ is the largest order such that:
@@
A spectral filter is an approximate inverse applied to the spectrum of a positive self-adjoint operator (e.g., the covariance/integral operator) via functional calculus: if a covariance operator $\mathsf{T}=\sum_{i}\sigma_{i}\langle\cdot,e_{i}\rangle e_{i}$, then ${\mathsf{g}}_{\lambda}(\mathsf{T})=\sum_{i}{\mathsf{g}}_{\lambda}(\sigma_{i})\langle\cdot,e_{i}\rangle e_{i}$. It attenuates directions with small eigenvalues $\sigma_{i}$ to stabilize the inversion, while the residual $1-\sigma{\mathsf{g}}_{\lambda}(\sigma)$ quantifies the bias introduced by regularization. The qualification $\nu$ is the largest order for which $r_{\lambda}(\sigma)\leq\lambda^{\nu}\sigma^{-\nu}$ uniformly. Typical values of $\nu$ are: Tikhonov with $\nu=1$; order-$t$ iterated Tikhonov with $\nu=t$; Landweber with $\nu\to\infty$ [11, Sec. 3.1]. The Lipschitz condition ensures stability of ${\mathsf{g}}_{\lambda}(\mathsf{T})$ to empirical perturbations.
@@ IV-D Source Condition
Target-kernel alignment is captured by an Hölder-type source condition [6] with exponent $r>0$; larger $r$ yields faster rates up to the filter qualification $\nu$. Conceptually, it measures how well $(\Phi_{\rho},f_{\rho})$ align with the vvRKHS prior. Formally, it requires membership in the range of a positive power of the integral operator $\mathsf{L}=\mathsf{I}\mathsf{I}^{*}$. Since SFL learns in a windowed flow space and then lifts guarantees to the vector field, we impose sources at both levels: a flow source driving the flow FS-HP bound and a vector-field source encoding physics, e.g., Hamiltonian passivity, incompressibility $\nabla\cdot f=0$, symmetry/equivariance. In practice, such structures can be enforced via physics-informed vvRKHS kernels, projection of the lifted estimator, or penalty terms [3, 1, 4]. For $r>0$ and $R_{\mathcal{V}}>0$, assume $f_{\rho}$ is the target vector field with the following structure defined through (4):
@@
where $\mathsf{L}_{\mathcal{W}}$ is defined in (5). This flow-level source is the one used in Lemma 3 to obtain the FS-HP rate $\ell^{-\frac{2r}{2r+\beta}}$ with $\beta=\max\{1,2\mu\}$.
@@ V Spectral Flow Learning Theory
This section develops SFL’s core machinery: variable-step windows yield well-posed labels via Koopman-lag forcing and state aggregation. We define the label map and vLMM residual, prove an observability inequality and a uniform local-truncation bound, and derive FS-HP excess-risk bounds for flows and for vector fields.
@@ V-A Label Map and vLMM Residual
A key aspect of a theory on learning a vector field from trajectory data is the construction of well-defined input-label pairs. In our framework, we leverage the Koopman-lag forcing operator $\mathsf{B}$ and state-aggregation operator $\mathsf{A}$ to define the population-level loss via the label map and vLMM residual. Consider a vLMM as defined in Definition 1. Fix the window and anchor it at $t_{k+M}$. Define the backward lags:
@@
Using the Koopman action $(\mathcal{K}^{t}f)(x)=f(\Phi^{t}(x))$, the right-hand side of (3) can be rewritten as:
$h_{k+M-1}\sum_{j=1}^{M}\beta_{jk}(\mathcal{K}^{-\tau_{M-j}}f)\big(x_{k+M}\big).$
@@
$(\mathsf{B}f)(x;h,\zeta)\coloneqq h\sum_{j=0}^{M}\beta_{j}(\zeta)(\mathcal{K}^{-\tau_{j}}f)(x),\quad\mathsf{B}:\mathcal{V}\to\mathcal{W},$ (6)
where $\tau_{j}\coloneqq\tau_{j}(\zeta)$ are the cumulative lags determined by $\zeta$ and the anchor step $h\coloneqq h_{k+M-1}$. Similarly, the state-aggregation operator is defined as:
$(\mathsf{A}\Phi)(x;h,\zeta)\coloneqq\sum_{j=1}^{M}\alpha_{j}(\zeta)\Phi^{-\tau_{j}}(x),\quad\mathsf{A}:\mathcal{W}^{\circ}\to\mathcal{W}.$
Now, we are ready to define the population-level label map and vLMM residual.
@@
The population-label map and vLMM residual are:
$\mathsf{Y}_{\alpha}(\Phi)\coloneqq\mathsf{J}\Phi+\mathsf{A}\Phi\in\mathcal{W},\ \Delta(f)\coloneqq\mathsf{Y}_{\alpha}(\Phi)-\mathsf{B}f\ \in\mathcal{W},$ (7)